50 ml sample of 0.0950 M acetic acid (ka: 1.8* 10-5) isbeing titrated with 0.106 M NaoH.
1- what is the PH at the midpoint of titration?
2- what is the PH equivalence point of the titration?
3-what is the PH and endpoint of titration/
Answer #1
Given data
ka = 1.8* 10-5
    pKa =4.74        
Volume of acetic acid, V1 = 50 mL
    Molarity of acetic acid M1 = 0.095M
Molarity of NaOH, M2 = 0.106 M
∴ The volume of NaOH are required to reachthe equivalence point, V2 = M1V1 / M2
                                                                                                                = 0.095 M * 50 mL / 0.106 M
                                                                                                                 =44.81 mL
At mid point of titartion number of moles of aceticacid = number of moles of sodium acetate
   ∴ pH = pKa + log[salt] /[acid]
[salt] /[acid] = 1 so  log[salt] /[acid] =0
    Hence pH = pKa
                    = 4.74
b) At equivalence point of the titration all the acidis converted into salt
∴ Concentration of salt = molarity of acid *volume of acid / total volume
                                      =0.095 M * 50 mL / ( 50 mL + 44.81 mL )
                                  C= 0.05 M
   Sodium acetate ca acts as a base
    The Kb of acetate ion = Kw/ Ka
                                     = 1.0*10-14 / 1.8*10-5
                                     = 5.56*10-10
The concentration of OH- ion at equivalence point =√Kb*C
                                                                              =(5.56*10-10 * 0.05)1/2
                                                                             = 5.27*10-6 M
                        ∴ pOH = - log [OH-]
                                      =- log( 5.27*10-6 M)
                                     = 5.278
                                pH = 14 - 5.278
                                      = 8.722
c) Similar to b)